OpenStudy (jackie):

shirt cost $20 sells for $33/jacket costs 60 and sells for 93. I know how to find the slope and set up the equation but then...it is said that if the markup policy of the store for items costing more than $10 is assumed to be linearly related, write the equation that expresses retail price R in terms of cost. ??

OpenStudy (jackie):

I am wondering if the $10 is to throw me off or I need to acutally set up a new equation because it asked for the new slope.

OpenStudy (anonymous):

key words are lineraly related. Therefore they are saying there is a realtionship between two items which cost more than ten dollars to make. Shirt costs 20$ and the jacket costs 60$. the line which is confusing you just states that the shirt and jacket are assumed to be realted or in other words on the same linear graph.

OpenStudy (anonymous):

93-33=60 / 60-20=40 final answear is 1.5 or 3/2

OpenStudy (anonymous):

the final equation would end up being sale price=1.5r+3